# The Open Journal of Astrophysics

### v1

@@ Abstract
To date, only two strongly lensed type Ia supernovae (SNIa) have been discovered with an isolated galaxy acting as the lens: iPTF16geu and SN Zwicky. The observed image fluxes for both lens systems were inconsistent with predictions from a smooth macro lens model. A potential explanation for the anomalous flux ratios is microlensing: additional (de)magnification caused by stars and other compact objects in the lens galaxy. In this work, we combine observations of iPTF16geu and SN Zwicky with simulated microlensing magnification maps, leveraging their standardizable candle properties to constrain the lens galaxy mass slope, $\eta$, and the fraction of dark compact objects, $f_{\rm dc}$. The resulting mass slopes are $\eta=1.70\pm 0.07$ for iPTF16geu and $\eta=1.81\pm 0.10$ for SN Zwicky. Our results indicate no evidence for a population of dark compact objects, placing upper limits at the $95\%$ confidence level of $f_{\rm dc}$ $<0.25$ for iPTF16geu and $f_{\rm dc}$ $<0.47$ for SN Zwicky. Assuming a constant fraction of dark compact objects for both lensed SNe, we obtain $f_{\rm dc}$ $<0.19$. These results highlight the potential of strongly lensed SNIa to probe the innermost parts of lens galaxies and learn about compact matter.
@@ 1 Introduction
Strong gravitational lensing occurs when light from a distant light source is deflected by the gravitational potential of a massive galaxy or galaxy cluster along the line of sight, resulting in multiple images of the event. When the light source is a transient event, such as a supernova (SN), we can measure the time delays between the appearance of the multiple images. Combined with a model for the gravitational potential of the lens galaxy and line-of-sight structures, this constrains the present-day expansion rate of the Universe, the Hubble constant (Refsdal, 1964; Treu & Marshall, 2016; Suyu et al., 2024; Birrer et al., 2024). Additionally, lensed SNe enable detailed studies of lensing galaxies and high-redshift SNe (Cano et al., 2018; Johansson et al., 2021; Dhawan et al., 2024).
@@
While strongly lensed SNe are very useful, they are also extremely rare and challenging to discover. To date, only eight multiply-imaged lensed SNe have been found. Six of them have been discovered behind galaxy clusters: ‘SN Refsdal’ (Kelly et al., 2015), ‘SN Requiem’ (Rodney et al., 2021), ‘AT2022riv’ (Kelly et al., 2022), ‘C22’ (Chen et al., 2022), ‘SN H0pe’ (Frye et al., 2024), and ‘SN Encore’ (Pierel et al., 2024). Additionally, two lensed SNe were lensed by isolated elliptical galaxies: ‘iPTF16geu’ (Goobar et al., 2017) and ‘SN Zwicky’ (Goobar et al., 2023), shown in Fig. 1. Both of these galaxy-scale lensed SNe were type Ia SNe (SNIa) that were recognised as being gravitationally lensed from their brightness relative to their redshifts. This could be inferred thanks to their standardizable-candle nature, and because they were bright enough to obtain a spectroscopic redshift. iPTF16geu was found with the intermediate Palomar Transient Factory (iPTF), a time-domain survey that ran from 2013 to 2017 (Kulkarni, 2013). From 2018 onwards, a new camera with a larger field of view was installed on the telescope and the survey continued as the Zwicky Transient Facility (ZTF; Bellm et al., 2019; Graham et al., 2019). After four years of ZTF observations, SN Zwicky was discovered. Simulation studies show that a handful more, unidentified, lensed SNe are expected to be present in the ZTF data below the spectroscopic follow-up limit (Sagués Carracedo et al., 2024). Archival studies have been conducted to search for these overlooked lensed SNe (Magee et al., 2023; Townsend et al., 2024), and in Townsend et al. (2024), two potential lensed SNe are identified. The number of lensed SN detections is expected to increase orders of magnitudes with the upcoming Vera Rubin Observatory (Wojtak et al., 2019; Goldstein et al., 2019; Oguri & Marshall, 2010; Sainz de Murieta et al., 2023; Arendse et al., 2024).
@@
iPTF16geu and SN Zwicky share several characteristics: they are both highly magnified ($67.8\pm 2.9$ and $23.7\pm 3.2$ times, respectively) and have short time delays of the order of a day (More et al., 2017; Dhawan et al., 2020; Goobar et al., 2023). Additionally, they are unusually compact lens systems: iPTF16geu has an Einstein radius of 0.29′′ and SN Zwicky of only 0.17′′. Compared to the population of known lensed galaxies and lensed quasars, both are outliers with small Einstein radii and faint lens galaxies, and they were only identified because they magnified SNIa standard candles (see Lemon et al., 2024, and Fig. 4 in Goobar et al., 2023). Another interesting shared property is that the flux ratios of their lensed images, after corrections for differential dust extinction in the lensing galaxy, are inconsistent with predicted flux ratios from a smooth macro lens model (More et al., 2017; Mörtsell et al., 2020; Goobar et al., 2023; Pierel et al., 2023). A potential explanation for the observed fluxes is additional (de)magnification from stars, stellar remnants and other compact objects in the lens galaxy; known as microlensing (Chang & Refsdal, 1979). Several papers have analysed SN Zwicky and found evidence that microlensing is needed to explain the observed flux ratios (Goobar et al., 2023; Pierel et al., 2023; Larison et al., 2024). Lensing at such small angular scales ($\sim 10^{-6}$′) results in unresolvable micro-images that sum to a total magnification which differs from the macro magnification of the lens galaxy (Young, 1981; Paczynski, 1986; Vernardos et al., 2024). This process perturbs the magnification of each image independently, resulting in flux ratios that are discrepant from the macro model predictions. Microlensing is often talked about in terms of complicating the use of lensed SNIa for time-delay cosmography (Dobler & Keeton, 2006; Yahalomi et al., 2017; Foxley-Marrable et al., 2018; Weisenbach et al., 2021, 2024). However, the combination of microlensing and lensed SNe Ia offers a unique opportunity: since SNe Ia are standardizable candles with predictable brightness for each lensed image, we are able to precisely detect the magnification contribution from microlensing.
@@
Microlensing allows us to probe the abundance of compact objects at the image positions, even if they are dark or too faint to observe. This makes microlensing a unique probe to distinguish between the smooth and compact matter components of the lens galaxy. By comparing the fraction of compact objects inferred from microlensing to the stellar mass fraction determined from photometry, we can determine if there are signs of an additional population of dark compact objects, such as primordial black holes (PBHs). PBHs are hypothetical black holes that formed in the early Universe through direct gravitational collapse (e.g. Khlopov, 2010; Carr et al., 2016; Escrivà et al., 2024). Since they are not formed through stellar collapse, their masses are not limited to the narrow mass range of stellar black holes. PBHs are candidates for part of the dark matter in the Universe, as well as potential seeds for supermassive black holes. In particular, gravitational wave observations from the Laser Interferometer Gravitational-Wave Observatory (LIGO) suggest that PBHs of intermediate mass ($10M_{\odot}<M<200M_{\odot}$) may constitute a substantial part of the dark matter in the Universe (Clesse & García-Bellido, 2015; Kashlinsky, 2016; Hawkins, 2020), which we could detect through microlensing observations.
@@
Microlensing flux ratio studies have been conducted extensively for strongly lensed quasars (e.g Chang & Refsdal, 1979; Wambsganss, 2006; Schmidt & Wambsganss, 2010; Vernardos & Tsagkatakis, 2019; Vernardos et al., 2024), of which several studies have constrained the abundance of compact objects through microlensing (e.g. Pooley et al., 2009, 2012; Schechter & Wambsganss, 2004; Schechter et al., 2014; Mediavilla et al., 2009, 2017; Esteban-Gutiérrez et al., 2020; Mediavilla & Jiménez-Vicente, 2024). Most of these studies conclude that the compact objects causing microlensing can be explained sufficiently by the normal stellar population, and the fraction of dark compact objects is constrained to be below $10\%$ for masses above $0.002M_{\odot}$ (Esteban-Gutiérrez et al., 2023; Mediavilla & Jiménez-Vicente, 2024). For galaxy-scale lensed SNe, the field of microlensing studies has only recently opened up, with iPTF16geu the first object that allowed such investigation. In Mörtsell et al. (2020), microlensing was included in the analysis of the image magnifications to obtain the lens galaxy mass slope, $\eta\sim 1.8$. Diego et al. (2022) studied the observed flux ratios and light curves of iPTF16geu and concluded that microlensing effects from stars suffice to explain the data. With the discovery of SN Zwicky, a second galaxy-scale lensed SN has become available to study microlensing. In this work, we perform a micro and macro lensing analysis of photometric observations of SN Zwicky and iPTF16geu, and develop a Bayesian model to constrain the lens galaxy mass slopes and compact objects fractions. By combining these with stellar mass fractions measured from photometry data, we estimate how much of the mass can be in the form of dark compact objects.
@@ 2 Observations
Here, we summarise the observations of iPTF16geu and SN Zwicky used in our analysis. More elaborate descriptions can be found in Goobar et al. (2017); Dhawan et al. (2020); Goobar et al. (2023); Pierel et al. (2023).
@@
SN iPTF16geu was discovered in 2016 with iPTF. Spectroscopic identification was carried out with Spectral Energy Distribution Machine (SEDM; Blagorodnova et al., 2018; Rigault et al., 2019) at the Palomar 60-inch telescope on 2 October 2016, and iPTF16geu was found to be consistent with a SNIa at $z=0.4$. For the purposes of this work, we focus on the high-resolution imaging data from the Keck II telescope on Mauna Kea and the Hubble Space Telescope (HST). The observations from the Keck telescope used in this work were conducted with the Near-IR Camera 2 (NIRC2; Fremling et al., 2022) in the $J$-band, using laser guided adaptive optics (LGS-AO). Details of the reduction can be found in Goobar et al. (2017); Dhawan et al. (2020). The HST data were taken with the Wide Field Camera 3, using the UV (WFC3/UVIS) and near-IR (WFC3/IR) channels. There were two epochs of observations in the $F390W$, $F475W$ filters and eight epochs in each of the $F625W$, $F814W$, $F110W$, and $F160W$ filters. After the SN faded, we obtained reference observations in each of the filters on 2018 November 10. The reference images both for the ground-based and HST data are a crucial benefit of lensed SNe, since we can model the lensed host galaxy without the presence of the point sources.
@@ 2.2 SN Zwicky
On August 2022, SN Zwicky was discovered with ZTF. The object was bright enough for an automatic classification with SEDM, where a spectrum taken on the 21st of August showed it to be a SNIa at a redshift of $z=0.35$, confirming a brightness several magnitudes brighter than a typical SNIa at that redshift. The multiple images of SN Zwicky were first resolved on 2022 September 15 with the Keck telescope in the $J$-band, using adaptive optics with the NIRC2 camera (Fremling et al., 2022). On September 21, the LensWatch collaboration [^1] https://www.lenswatch.org/home observed SN Zwicky using HST with the optical filters $F475W$, $F625W$ and $F814W$ (WFC3/UVIS Camera) and the NIR $F160W$ filter (WFC3/IR) (Pierel et al., 2022, 2023). To allow for further studies of the lens and host galaxies, reference images after SN Zwicky faded were obtained with HST in the same filters (Larison et al., 2024) and also with the James Webb Space Telescope (JWST) [^2] Program ID: GO-2905 in the filters $F115W$, $F150W$, $F356W$, and $F444W$, as shown in a composite image in Fig. 1.
@@ 3.1 Mass and light model definitions
where $I_{\rm e}$ is the intensity at the half-light radius $R_{\rm e}$, $b_{n}=1.9992n-0.3271$ (Birrer & Amara, 2018) and
@@ 3.2 Fitting methodology
Using the open-source lens modelling software Lenstronomy [^3] https://lenstronomy.readthedocs.io/en/latest/ (Birrer & Amara, 2018; Birrer et al., 2021), we simultaneously reconstructed the lens mass model, SN image positions, and the light distributions of the lens and host galaxy. For both iPTF16geu and SN Zwicky, we perform lens modelling for several observations from different telescopes and in different filters, each with characteristics that require slightly different approaches to the modelling. In this section, we outline our overarching modelling approach which applies to most observations, and we highlight deviations from this for specific cases in their respective sections.
@@
We iterate three times over fitting preliminary models with particle swarm optimization (PSO). For each of the three runs of PSO, we use 100, 100, and 10 particles and 100, 100, and 10 iterations, respectively. For the first run, we initialise the particles according to wide Gaussian priors around the expected parameter values. For the second and third runs, we reduce the standard deviations of the particle initial positions to 10% and 1% of their original values, respectively, around the best fit parameter values from the previous PSO iteration. We also iterate over the PSF model, using a built-in feature in Lenstronomy that minimises the residuals between the observed and reconstructed image around the SN positions (Shajib et al., 2019). Once a suitable PSF is obtained, we run a Markov chain Monte Carlo (MCMC) sampler to obtain parameter estimates with uncertainties. In particular, we use an affine invariant MCMC ensemble sampler from python package emcee (Foreman-Mackey et al., 2013; Goodman & Weare, 2010). We typically use 200 walkers with 500 burn-in steps that we discard and 1500 steps from which we obtain the final posterior on the model parameters. Because we have already used PSO to bring the fitter close to the posterior maximum, we initialize the walkers with a standard deviation of 2% of that specified in the Gaussian priors.

### v2

@@ Abstract
To date, only two strongly lensed type Ia supernovae (SNIa) have been discovered with an isolated galaxy acting as the lens: iPTF16geu and SN Zwicky. The observed image fluxes for both lens systems were inconsistent with predictions from a smooth macro lens model. A potential explanation for the anomalous flux ratios is microlensing: additional (de)magnification caused by stars and other compact objects in the lens galaxy. In this work, we combine observations of iPTF16geu and SN Zwicky with simulated microlensing magnification maps, leveraging their standardizable candle properties to constrain the lens galaxy mass slope, $\eta$, and the fraction of dark compact objects, $f_{\rm dc}$. The resulting mass slopes are $\eta=1.70\pm 0.07$ for iPTF16geu and $\eta=1.81\pm 0.10$ for SN Zwicky. Our results indicate no evidence for a population of dark compact objects, placing upper limits at the $95\%$ confidence level of $f_{\rm dc}$ <0.25<0.25 for iPTF16geu and $f_{\rm dc}$ <0.47<0.47 for SN Zwicky (for compact objects with masses above $0.02M_{\odot}$). Assuming a constant fraction of dark compact objects for both lensed SNe, we obtain $f_{\rm dc}$ <0.19<0.19. These results highlight the potential of strongly lensed SNIa to probe the innermost parts of lens galaxies and learn about compact matter.
@@ 1 Introduction
Strong gravitational lensing occurs when light from a distant light source is deflected by the gravitational potential of a massive galaxy or galaxy cluster along the line of sight, resulting in multiple images of the event. When the light source is a transient event, such as a supernova (SN), we can measure the time delays between the appearance of the multiple images. Combined with a model for the gravitational potential of the lens galaxy and line-of-sight structures, this constrains the present-day expansion rate of the Universe, the Hubble constant $H_{0}$ (refsdal1964hubble; treu2016cosmography; Suyu2024; Birrer2024). Additionally, lensed SNe enable detailed studies of lensing galaxies and high-redshift SNe (cano2018spectroscopic; Johansson2021; Dhawan2024).
@@
While strongly lensed SNe are very useful, they are also extremely rare and challenging to discover. To date, only eight multiply-imaged lensed SNe have been found. Six of them have been discovered behind galaxy clusters: ‘SN Refsdal’ (kelly2015multiple), ‘SN Requiem’ (rodney2021gravitationally), ‘AT2022riv’ (Kelly2022riv), ‘C22’ (Chen2022_glSNAbell), ‘SN H0pe’ (Frye2024), and ‘SN Encore’ (Pierel2023_Encore). Additionally, two lensed SNe were lensed by isolated elliptical galaxies: ‘iPTF16geu’ (goobar2017iptf16geu) and ‘SN Zwicky’ (Goobar2023_SNZwicky), shown in Fig. 1. Both of these galaxy-scale lensed SNe were type Ia SNe (SNIa) that were recognised as being gravitationally lensed from their brightness relative to their redshifts. This could be inferred thanks to their standardizable-candle nature, and because they were bright enough to obtain a spectroscopic redshift. iPTF16geu was found with the intermediate Palomar Transient Factory (iPTF), a time-domain survey that ran from 2013 to 2017 (Kulkarni2013_iPTF). From 2018 onwards, a new camera with a larger field of view was installed on the telescope and the survey continued as the Zwicky Transient Facility (ZTF; bellm2019ZTF; Graham2019_ZTF). After four years of ZTF observations, SN Zwicky was discovered. Simulation studies show that a handful more, unidentified, lensed SNe are expected to be present in the ZTF data below the spectroscopic follow-up limit (Sagues2024). Archival studies have been conducted to search for these overlooked lensed SNe (Magee2023; Townsend2024), and in Townsend2024, two potential lensed SNe are identified. The number of lensed SN detections is expected to increase orders of magnitudes with the upcoming Vera Rubin Observatory (wojtak2019magnified; goldstein2019rates; oguri2010gravitationally; SainzdeMurieta2023; Arendse2024).
@@
iPTF16geu and SN Zwicky share several characteristics: they are both highly magnified ($67.8\pm 2.9$ and $23.7\pm 3.2$ times, respectively) and have short time delays of the order of a day (more2017interpreting; Dhawan2020_16geu; Goobar2023_SNZwicky). Additionally, they are unusually compact lens systems: iPTF16geu has an Einstein radius of 0.29′′ and SN Zwicky of only 0.17′′. Compared to the population of known lensed galaxies and lensed quasars, both are outliers with small Einstein radii and faint lens galaxies, and they were only identified because they magnified SNIa standard candles (see Lemon2024, and Fig. 4 in Goobar2023_SNZwicky). Another interesting shared property is that the flux ratios of their lensed images, after corrections for differential dust extinction in the lensing galaxy, are inconsistent with predicted flux ratios from a smooth macro lens model (more2017interpreting; Mortsell2020; Goobar2023_SNZwicky; Pierel2023). A potential explanation for the observed fluxes is additional (de)magnification from stars, stellar remnants and other compact objects in the lens galaxy; known as microlensing (Chang1979). Several papers have analysed SN Zwicky and found evidence that microlensing is needed to explain the observed flux ratios (Goobar2023_SNZwicky; Pierel2023; Larison2024). Lensing at such small angular scales ($\sim 10^{-6}$ ′′) results in unresolvable micro-images that sum to a total magnification which differs from the macro magnification of the lens galaxy (Young1981; Paczynski1986; Vernardos2024). This process perturbs the magnification of each image independently, resulting in flux ratios that are discrepant from the macro model predictions. Microlensing is often talked about in terms of complicating the use of lensed SNIa for time-delay cosmography (dobler2006microlensing; Yahalomi2017; foxley2018standardization; Weisenbach2021; Weisenbach_2024). However, the combination of microlensing and lensed SNe Ia offers a unique opportunity: since SNe Ia are standardizable candles with predictable brightness for each lensed image, we are able to precisely detect the magnification contribution from microlensing.
@@
Microlensing allows us to probe the abundance of compact objects at the image positions, even if they are dark or too faint to observe. This makes microlensing a unique probe to distinguish between the smooth and compact matter components of the lens galaxy. By comparing the fraction of compact objects inferred from microlensing to the stellar mass fraction determined from photometry, we can determine if there are signs of an additional population of dark compact objects, such as primordial black holes (PBHs). PBHs are hypothetical black holes that formed in the early Universe through direct gravitational collapse (e.g. Khlopov2010; Carr2016; Escriva2022). Since they are not formed through stellar collapse, their masses are not limited to the narrow mass range of stellar black holes. PBHs are candidates for part of the dark matter in the Universe, as well as potential seeds for supermassive black holes. In particular, gravitational wave observations from the Laser Interferometer Gravitational-Wave Observatory (LIGO) suggest that PBHs of intermediate mass ($10M_{\odot}<M<200M_{\odot}$) may constitute a substantial part of the dark matter in the Universe (Clesse2015; Kashlinsky2016; Hawkins2020), which we could detect through microlensing observations.
@@
Microlensing flux ratio studies have been conducted extensively for strongly lensed quasars (e.g Chang1979; Wambsganss2006; Schmidt2010; Vernardos2019; Vernardos2024), of which several studies have constrained the abundance of compact objects through microlensing (e.g. Pooley2009; Pooley2012; Schechter2004; Schechter2014; Mediavilla2009; Mediavilla2017; Esteban2020; Medavilla2024). Most of these studies conclude that the compact objects causing microlensing can be explained sufficiently by the normal stellar population, and the fraction of dark compact objects is constrained to be below $10\%$ for masses above $0.002M_{\odot}$ (Gutierrez2023; Medavilla2024). For galaxy-scale lensed SNe, the field of microlensing studies has only recently opened up, with iPTF16geu the first object that allowed such investigation. In Mortsell2020, microlensing was included in the analysis of the image magnifications to obtain the lens galaxy mass slope, $\eta\sim 1.8$. Diego2022 studied the observed flux ratios and light curves of iPTF16geu and concluded that microlensing effects from stars suffice to explain the data. With the discovery of SN Zwicky, a second galaxy-scale lensed SN has become available to study microlensing. In this work, we perform a micro and macro lensing analysis of photometric observations of SN Zwicky and iPTF16geu, and develop a Bayesian model to constrain the lens galaxy mass slopes and compact objects fractions. By combining these with stellar mass fractions measured from photometry data, we estimate how much of the mass can be in the form of dark compact objects.
@@ 2 Observations
Here, we summarise the observations of iPTF16geu and SN Zwicky used in our analysis. More elaborate descriptions can be found in goobar2017iptf16geu; Dhawan2020_16geu; Goobar2023_SNZwicky; Pierel2023.
@@
SN iPTF16geu was discovered in 2016 with iPTF. Spectroscopic identification was carried out with Spectral Energy Distribution Machine (SEDM; Blago2018; Rigault2019) at the Palomar 60-inch telescope on 2 October 2016, and iPTF16geu was found to be consistent with a SNIa at $z=0.4$. For the purposes of this work, we focus on the high-resolution imaging data from the Keck II telescope on Mauna Kea and the Hubble Space Telescope (HST). The observations from the Keck telescope used in this work were conducted with the Near-IR Camera 2 (NIRC2; 2022Fremling) in the $J$-band, using laser guided adaptive optics (LGS-AO). Details of the reduction can be found in goobar2017iptf16geu; Dhawan2020_16geu. The HST data were taken with the Wide Field Camera 3, using the UV (WFC3/UVIS) and near-IR (WFC3/IR) channels. There were two epochs of observations in the $F390W$, $F475W$ filters and eight epochs in each of the $F625W$, $F814W$, $F110W$, and $F160W$ filters. After the SN faded, we obtained reference observations in each of the filters on 2018 November 10. The reference images both for the ground-based and HST data are a crucial benefit of lensed SNe, since we can model the lensed host galaxy without the presence of the point sources.
@@ 2.2 SN Zwicky
On August 2022, SN Zwicky was discovered with ZTF. The object was bright enough for an automatic classification with SEDM, where a spectrum taken on the 21st of August showed it to be a SNIa at a redshift of $z=0.35$, confirming a brightness several magnitudes brighter than a typical SNIa at that redshift. The multiple images of SN Zwicky were first resolved on 2022 September 15 with the Keck telescope in the $J$-band, using adaptive optics with the NIRC2 camera (2022Fremling). On September 21, the LensWatch collaboration [^1] https://www.lenswatch.org/home observed SN Zwicky using HST with the optical filters $F475W$, $F625W$ and $F814W$ (WFC3/UVIS Camera) and the NIR $F160W$ filter (WFC3/IR) (2022Pierel; Pierel2023). To allow for further studies of the lens and host galaxies, reference images after SN Zwicky faded were obtained with HST in the same filters (Larison2024) and also with the James Webb Space Telescope (JWST) [^2] Program ID: GO-2905 in the filters $F115W$, $F150W$, $F356W$, and $F444W$, as shown in a composite image in Fig. 1.
@@ 3.1 Mass and light model definitions
where $I_{\rm e}$ is the intensity at the half-light radius $R_{\rm e}$, $b_{n}=1.9992n-0.3271$ (birrer2018lenstronomy) and
@@
3.2 Testing angular complexity in the mass model
To investigate whether deviations from a smooth elliptical mass distribution can account for the observed flux anomalies in the SN images, we tested the inclusion of multipole perturbations in the lens model. Recent work by Cohen2024 has shown that higher-order angular structure in the lens galaxy can also reproduce flux-ratio anomalies, potentially providing an alternative explanation to microlensing. Specifically, we considered both circular and elliptical fourth-order multipole perturbations to the convergence, following the formalism of Xu2013 as implemented in the open-source lens modelling software Lenstronomy [^3] https://lenstronomy.readthedocs.io/en/latest/ (birrer2018lenstronomy; Birrer2021LenstronomyII). The fourth-order multipole was added as an additional mass component with two free parameters: the amplitude $a_{4}$ and the phase $\phi_{4}$. These multipole terms allow for departures from purely elliptical isodensity contours, introducing boxy or disky shapes. We applied this extended mass model to Keck $J$-band observations of both SN Zwicky and iPTF16geu, fixing the mass slope $\eta$ to our best-fit values. The flux ratios of the SN images were included in the likelihood function to assess whether the lens model could reproduce them.
We briefly describe our main findings here. Our results for iPTF16geu yield a poor fit, with a high reduced $\chi^{2}$ value. This demonstrates that even with the included angular complexity in the mass distribution, the lens model fails to reproduce the observed flux ratios. For SN Zwicky, the multipole model matches the observed flux ratios more closely, but the resulting mass distribution is highly irregular and unphysical, with best-fit multipole amplitudes ($a_{4}=-0.192^{+0.007}_{-0.005}$ for circular multipoles and $a_{4}=-0.103^{+0.006}_{-0.005}$ for elliptical multipoles) inconsistent with realistic galaxy profiles (which are often of the order $10^{-2}$, see e.g. Hao2006). These unphysical multipole amplitudes are likely a consequence of the absence of a lensed host galaxy in the SN Zwicky observations, which would otherwise have constrained the model to more realistic values, as is the case for iPTF16geu. We conclude that angular complexity does not provide a plausible physical explanation for the observed flux ratios in either lensed SN system. These results strengthen the case that microlensing is required to explain the observed flux anomalies.
When the flux ratio constraints are excluded, adding multipole perturbations to the lens model decreases the reduced $\chi^{2}$ by only 0.008 for iPTF16geu and provides no improvement for SN Zwicky. In our subsequent analysis, we do not include angular complexity and proceed with the mass model described in Section 3.1.
@@ 3.3 Fitting methodology
Using Lenstronomy, we simultaneously reconstruct the lens mass model, SN image positions, and the light distributions of the lens and host galaxy. For both iPTF16geu and SN Zwicky, we perform lens modelling for several observations from different telescopes and in different filters, each with characteristics that require slightly different approaches to the modelling. In this section, we outline our overarching modelling approach which applies to most observations, and we highlight deviations from this for specific cases in their respective sections.
@@
We iterate three times over fitting preliminary models with particle swarm optimization (PSO). For each of the three runs of PSO, we use 100, 100, and 10 particles and 100, 100, and 10 iterations, respectively. For the first run, we initialise the particles according to wide Gaussian priors around the expected parameter values. For the second and third runs, we reduce the standard deviations of the particle initial positions to 10% and 1% of their original values, respectively, around the best fit parameter values from the previous PSO iteration. We also iterate over the PSF model, using a built-in feature in Lenstronomy that minimises the residuals between the observed and reconstructed image around the SN positions (Shajib2019). Once a suitable PSF is obtained, we run a Markov chain Monte Carlo (MCMC) sampler to obtain parameter estimates with uncertainties. In particular, we use an affine invariant MCMC ensemble sampler from python package emcee (ForemanMackey2013; Goodman2010). We typically use 200 walkers with 500 burn-in steps that we discard and 1500 steps from which we obtain the final posterior on the model parameters. Because we have already used PSO to bring the fitter close to the posterior maximum, we initialize the walkers with a standard deviation of 2% of that specified in the Gaussian priors.